### v1

# A Comprehensive Benchmark for Electrocardiogram Time-Series

Abstract.
Electrocardiogram (ECG), a key bioelectrical time-series signal, is crucial for assessing cardiac health and diagnosing various diseases. Given its time-series format, ECG data is often incorporated into pre-training datasets for large-scale time-series model training. However, existing studies often overlook its unique characteristics and specialized downstream applications, which differ significantly from other time-series data, leading to an incomplete understanding of its properties. In this paper, we present an in-depth investigation of ECG signals and establish a comprehensive benchmark, which includes (1) categorizing its downstream applications into four distinct evaluation tasks, (2) identifying limitations in traditional evaluation metrics for ECG analysis, and introducing a novel metric; (3) benchmarking state-of-the-art time-series models and proposing a new architecture. Extensive experiments demonstrate that our proposed benchmark is comprehensive and robust. The results validate the effectiveness of the proposed metric and model architecture, which establish a solid foundation for advancing research in ECG signal analysis.
Electrocardiogram, Timer-Series Model, Datasets and Benchmark, Biological Application
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYconference: Proceedings of the 33th ACM International Conference on Multimedia; October 27-31, 2025; Dublin, lrelandsubmissionid: 332ccs: Computing methodologiesccs: Computing methodologies Artificial intelligence
1. Introduction
Figure 1. Illustrations of our proposed ECG benchmark. (a) Overview of four ECG evaluation tasks, including classification, detection, forecasting, and generation, across various ECG applications. (b) Comparison between traditional ECG metric MSE and our proposed metric, FFD. The results are computed using the forecasting results of Medformer (Wang et al., 2024) and Timer (Liu et al., 2024) on the NFE dataset (Silva et al., 2013a); more details are provided in Section 4.3. (c) Performance comparison of ECG models, including traditional methods, large time series models, and our proposed Patch Step-by-Step Model (PSSM).
[…]
The electrocardiogram (ECG) is a time-series representation that records the heart’s electrical activity, describing the potential variations of myocardial cells during the depolarization and repolarization processes (Bonow et al., 2011). It is collected through electrodes placed on the chest and limbs (Macfarlane et al., 2010). In traditional analysis, the main components of an ECG are the P-wave (atrial depolarization), QRS complex (ventricular depolarization), and T-wave (ventricular repolarization) (Kligfield et al., 2007), as shown in Figure 1(a). Abnormalities in these waveforms and their temporal domains can be used to monitor cardiac health, such as arrhythmia classification (Fuster et al., 2006) and early detection of cardiovascular events (Reichlin et al., 2009), and to diagnose cardiac diseases, such as myocardial ischemia and electrolyte imbalances (Bonow et al., 2011). Researchers usually collect ECG data for their research objective, employing traditional methods such as time-frequency analysis (Van Geijn et al., 1980; Goldberger et al., 2000). However, with the advent of deep learning and large-scale training, the artisanal paradigm has become obsolete due to its limitations, such as insufficient data (Miotto et al., 2018) and diverse research goals (Rajpurkar et al., 2017). Therefore, integrating and standardizing ECG data and its downstream objectives into a unified benchmark has become an urgent requirement for advancing this field.
Inspired by the success of large language models, pre-training large time-series models (LTMs) has become a mainstream trend in time-series modeling (Liu et al., 2024; Gao et al., 2024). As a time-series data type, ECG is often integrated with other time-series datasets, such as meteorological and financial data, to pre-train LTMs (Ma et al., 2024). However, this integration overlooks the unique characteristics of ECG, leading to potential misunderstandings. Compared to other time-series data, ECG has the following features: (1) It exhibits quasi-periodic behavior, which differs from other time-series with no periodicity (e.g., financial data) or stable periodicity (e.g., weather data). This is because the periodicity of ECG is influenced by physiological factors, such as heart rate and myocardial ischemia (Wagner and Strauss, 2014; Chakraborty et al., 2018); (2) ECG is applied in a wide range of complex applications, such as disease diagnosis based on classification and maternal-fetal waveform prediction based on generations. These tasks differ from the primarily predictive goals of other time-series data, like stock price forecasting (Liu et al., 2024; Zhou et al., 2022). Therefore, it is necessary to construct a specialized benchmark for ECG to validate existing methods and guide the design of new ones.
In this paper, we propose a comprehensive benchmark to standardize ECG learning and evaluation, comprising the following three parts:
(1) Comprehensive ECG evaluations. As shown in Figure 1(a), we define a set of evaluations designed for the broad range of medical applications of ECG, which include four downstream tasks: (a) Classification for disease diagnosis and event prediction (Hannun et al., 2019; Ribeiro et al., 2020), (b) Detection for key waveform (e.g., P-wave) localization  (Murray et al., 1980; Kumar et al., 2019), (c) Forecasting for ECG dynamics prediction (Prakarsha and Sharma, 2022), (d) Generation for maternal-fetal ECG separation (Sarkar and Etemad, 2021). These tasks provide a comprehensive framework for evaluation, which can fully assess model performance and ensure that they are rigorously tested and can be effectively adapted to various ECG applications.
(2) Innovative ECG Metric. The Mean Squared Error (MSE), commonly used for time-series evaluations, is often unsuitable for assessing ECG and may lead to misleading results. For example, as shown in Figure 1(b), generated ECGs that preserve clinically valid morphological patterns but exhibit minor temporal shifts can lead to large MSE, even higher than those of a meaningless flat line prediction. This phenomenon arises from the quasi-periodic nature and extreme values in ECG signals, highlighting the need for more suitable evaluation metrics. Inspired by image quality metrics (Heusel et al., 2017), we introduce the Feature-based Fréchet Distance (FFD) to assess the quality of generated ECG signals. FFD calculates the distance between real and generated ECGs by comparing their latent feature distributions, providing a more semantically meaningful measure of generated ECG quality. Additionally, this approach offers improved robustness by better aligning the evaluation with clinical interpretations of ECG morphology.
(3) ECG-Specific Model. Traditional time-series models may be suboptimal for ECG feature extraction due to its unique characteristics. Inspired by the cardiac conduction system (Trayanova et al., 2023), we propose the Patch Step-by-Step Model (PSSM), an encoder-centric hierarchical architecture that adaptively captures cross-scale features from localized waveform segments to global rhythm patterns through iterative signal patching. Extensive experiments demonstrate the effectiveness of our proposed model on the benchmark and metric we have introduced.
Our main contributions can be summarized as follows:
  1. (1)
    We highlight the unique characteristics of ECG signals compared to other time-series data and analyze the limitations of traditional methods and large time-series models in processing ECG. This underscores the need for a comprehensive ECG benchmark.
  2. (2)
    We propose an ECG benchmark comprising: (a) comprehensive multi-task ECG evaluations, integrating classification, detection, forecasting, and generation to represent the underlying applications of ECG; (b) a novel metric, Feature-based Fréchet Distance (FFD), as a supplement to MSE, to assess semantic fidelity of ECG signals; (c) a novel architecture, Patch Step-by-Step Model (PSSM), specifically tailored for ECG, which hierarchically encodes ECG signals through adaptive patching.
  3. (3)
    We conduct extensive experiments to validate the rationale of our ECG benchmark, the robustness of FFD, as well as the effectiveness of our proposed PSSM. The experimental results demonstrate that the proposed benchmark and metric are necessary, and PSSM achieves state-of-the-art performance, providing a solid foundation for advancing this field.
Code: https://github.com/ZhijiangTang/ECG-Benchmark
2. Related Works
ECG Analysis Methods. Current ECG analysis can be categorized into two main camps: (1) Traditional signal processing-based methods. For example, researchers used Fourier analysis and wavelet transform for QRS complex detection (Murray et al., 1980; Kumar et al., 2019) and applied RR interval variability for arrhythmia classification (Tsipouras et al., 2002). (2) Deep learning-based methods, where researchers applied deep models to significantly improve the performance of ECG tasks. For example, they used Tranformer-based frameworks for disease classification (El-Ghaish and Eldele, 2024; Miotto et al., 2018; Liu et al., 2021), CNN-based frameworks for waveforms detection (Peimankar and Puthusserypady, 2021), and GAN-based frameworks for ECG generation (Sarkar and Etemad, 2021; Mohebbian et al., 2021). However, these methods are designed and validated for specific ECG tasks and cannot fully capture the complex characteristics of ECG signals. In this paper, we propose a comprehensive ECG benchmark that thoroughly evaluates the ability of time-series models to address a wide variety of ECG applications.
Figure 2. Illustrations of the architecture of our Patch Step-by-Step Model (PSSM). (a) Components of the ConvBlock. (b) Patching operation in the encoder, where the tokens are obtained by averaging the corresponding tokens from the previous layer. (c) Unpatching operation in the decoder, where tokens are generated by the weighted splitting of the corresponding tokens from the last layer. (d) Projection heads for the four ECG evaluation tasks.
[…]
Time Series Foundational Models. Time-series analysis has evolved from classical statistical methods like autoregressive integrated moving average model (Box et al., 2015), which models temporal dependencies through autoregression and moving averages, to deep learning approaches such as recurrent neural networks (Rumelhart et al., 1986), long short-term memory model (Hochreiter, 1997), and transformers (Vaswani, 2017), which capture long-range dependencies with recurrent or self-attention mechanisms. Recent advancements have focused on improving transformer-based architectures for time-series tasks: Informer (Zhou et al., 2021) introduced linear complexity sparse attention mechanisms for long sequences, while Medformer (Wang et al., 2024) tailored transformers for medical time-series featuring via multi-scale data fusion.
Large Time-Series Models (LTMs). LTMs can be categorized into two main approaches: (1) Direct application of large language models (LLMs) to time-series modelling. For example, OneFitsAll (Zhou et al., 2023) leveraged GPT-2 for time-series forecasting, and Time-LLM (Jin et al., 2023) enhanced zero-shot generalization of LLMs through prompt engineering. (2) Pre-training models on large-scale time-series datasets (Dooley et al., 2024; Ekambaram et al., 2023, 2024). For example, Timer (Liu et al., 2024) pre-trained a Transformer decoder using next-token prediction for time-series forecasting, and UniTS (Gao et al., 2024) extended Transformer encoder with multi-task heads and jointly pre-trained on multimodal time-series data using a masked reconstruction objective. Furthermore, ECGPT (Davies et al., 2024) pre-trained a Transformer decoder on ECG data, while its small pre-training dataset and singular focus limited its ability to address the diverse downstream applications of ECG. In this paper, we highlight the unique characteristics of ECG compared to other time-series data and argue that LTMs may not necessarily be suitable for various ECG applications. Our extensive experiments demonstrate our benchmark’s rationality, our metric’s robustness, and our model’s effectiveness, offering new insights for the development of ECG analysis.
3. Method
3.1. Preliminary: Time-series Models
The training of time-series models mimics the training of large language models, leveraging a self-supervised framework based on the next token prediction (Liu et al., 2024; Gao et al., 2024). Given a time-series dataset $\mathcal{D}\!=\!\{\bm{x}\}$, each sample $\bm{x}$ with time length $t$ is initially patched into a sequence of $n$ time tokens, i.e., $\bm{x}\!=\!\{s_{1},s_{2},\dots,s_{n}\}$, where each token spans a time length of $m=t/n$. Then, the loss function for training time-series models can be written as:
(1) $\mathcal{L}_{\text{NTP}}=\frac{1}{n\!-\!1}\sum_{i=1}^{n-1}\frac{1}{m}\left\|s_{i+1}-e_{i}w\right\|^{2},$
where $s_{i+1}\!\in\!\mathbb{R}^{1\times m}$ is the ground truth for the $(i\!+\!1)$-th time-series tokens with $m$ time length, $e_{i}\!\in\!\mathbb{R}^{1\times d}$ denotes the features of tokens at $s_{i}$ extracted by the time-series model, $d$ denotes the hidden dimension, and $w\in\mathbb{R}^{d\times m}$ is a linear layer aligning the hidden features with the prediction space.
As a type of time-series signal, ECG can also be trained using Eq. (1). Given the diverse downstream applications and unique characteristics of ECG, we next introduce our benchmark with four evaluations, a new metric for ECG, and our proposed ECG models.
3.2. ECG Evaluations
Classification. ECG can be used to classify diseases, such as atrial fibrillation and hyperkalemia classification, based on distinct ECG patterns from different patients (Macfarlane et al., 2010). We thus formulate the classification task as an ECG evaluation, where the input is an individual’s ECG, i.e., $\bm{x}$, and the output is its corresponding class label $y$. The loss can be written as:
(2) $\mathcal{L}_{\text{cls}}=y\log({\frac{\exp(\bar{\bm{e}}w)}{\sum_{k}\exp(\bar{\bm{e}}w)_{k}}}),$
where $y$ denotes the one-hot class label, $\exp$ denotes the exponential function, $\bar{\bm{e}}\!\in\!\mathbb{R}^{1\times d}$ is the averaged feature of $\bm{e}\!=\!\{e_{1},e_{2},\ldots,e_{n}\}$, extracted from $\bm{x}\!=\!\{s_{1},s_{2},\dots,s_{n}\}$ by models, $d$ is the feature dimension, $w\!\in\!\mathbb{R}^{d\times c}$ is a linear layer, $c$ is the number of classes, and $k$ indexes all classes.
Detection. ECG can be used to calculate cardiovascular metrics (e.g., heart rate variability) by detecting key waveform positions (Peimankar and Puthusserypady, 2021). Thus, we formulate detection as an evaluation for ECG, where the given input is $\bm{x}$. The output is the position of key waveforms, annotated with a binary waveform label $\bm{y}$, which spans the same duration $t$ as $\bm{x}$. The loss for detection can be written as:
(3) $\mathcal{L}_{\text{det}}=\bm{y}\log\left(\sigma(\bm{e}w)\right)+(1\!-\!\bm{y})\log\left(1\!-\!\sigma(\bm{e}w)\right),$
where $\bm{y}\!\in\!\mathbb{R}^{1\times t}$ is the binary label, $\sigma$ is the sigmoid function, $\bm{e}\!\in\!\mathbb{R}^{n\times d}$ is the extracted feature for $\bm{x}$, and $w\!\in\!\mathbb{R}^{d\times m}$ is the linear projector. Note that, to align with the dimensions of $\bm{y}$, a reshape operation is applied to $\bm{e}w\in\mathbb{R}^{n\times m}$ to convert its dimension into $\mathbb{R}^{nm\times 1}$ (i.e., $\mathbb{R}^{t\times 1}$), which is omitted in the equation for simplicity.
Forecasting. ECG enables early risk alerts and facilitates personalized treatment by predicting its changes (Reichlin et al., 2009). We thus define the forecasting task, where the input is an ECG time-series $\bm{x}\!=\!\{s_{1},s_{2},\dots,s_{n}\}$, and the output is the subsequent ECG $\bm{x}^{\prime}\!=\!\{s_{n+1},s_{n+2},\dots,s_{n+n^{\prime}}\}$, and the loss is similar to Eq. (1):
(4) $\mathcal{L}_{\text{forecast}}=\frac{1}{n^{\prime}\!-\!1}\sum_{i=n}^{n+n^{\prime}\!-\!1}\frac{1}{m}\left\|s_{i+1}-e_{i}w\right\|^{2},$
where the notations have the same meanings as in Eq. (1).
Generation. In clinical practice, acquired ECGs often contain noise (Chatterjee et al., 2020) or require invasive acquisition (e.g., fetal ECG (Silva et al., 2013b)). Therefore, we define the generation task, where the input is an easily accessible ECG $\bm{x}$ with time length $t$, and the output is a corresponding ECG $\bm{y}$ aligned with $\bm{x}$. The loss is defined as:
(5) $\mathcal{L}_{\text{gen}}=\frac{1}{t}\sum_{i=1}^{t}\left\|y_{i}-e_{i}w\right\|^{2},$
where $y_{i}$ is the ground truth ECG at time $i$ of $\bm{y}$, $e_{i}\in\mathbb{R}^{1\times d}$ is the features of $\bm{x}$ at time $i$ extracted by the model, and $w\in\mathbb{R}^{d\times 1}$ is a linear projector.
3.3. ECG Metric: Feature-based Fréchet Distance
Mean Squared Error (MSE) is widely applied in temporal forecasting evaluations (Liu et al., 2024; Gao et al., 2024; Zhou et al., 2021; Nie et al., 2022) due to its simplicity and efficiency. However, it exhibits two critical limitations for ECG assessment: (1) it fails to capture the semantic fidelity of ECG (e.g., the quasi-periodic characteristics), and (2) it is sensitive to extreme values (e.g., the R-wave). Figure 1(b) shows that generated ECG preserving clinically valid patterns yields higher MSE than meaningless flat line predictions when minor temporal shifts occur. Inspired by image quality metrics (Heusel et al., 2017), we propose a new metric, Feature-based Fréchet Distance (FFD), as a complementary of MSE.
Consider a mapping $f\!:\!\mathbb{R}^{t}\!\rightarrow\!\mathbb{R}^{k}$ that projects ECG signals $\bm{x}$ into normally distributed features $e\sim\mathcal{N}(\mu,\Sigma)$. We quantify the differences between generated and real ECG Feature distributions using the Fréchet Distance:
[…]
where Tr is the trace operator, $\mu$ and $\Sigma$ denote the mean vector and covariance matrix of real ECG features, while $\hat{\mu}$ and $\hat{\Sigma}$ denote those of generated ECG features.
To calculate Eq. (6), empirically, $\mu,\Sigma$ could be estimated by $\tilde{\mu},\tilde{\Sigma}$ as follows:
$\tilde{\mu}=\frac{1}{N}\sum_{i=1}^{N}\bm{e}_{i},\quad\tilde{\Sigma}=\frac{1}{N-1}\sum_{i=1}^{N}(\bm{e}_{i}-\tilde{\mu})^{\top}(\bm{e}_{i}-\tilde{\mu})$,
where $N$ denotes the number of real ECG samples, $\bm{e}_{i}$ represents the feature of $\bm{x}_{i}$ mapped by $f$, which could be implemented as a standard time-series feature extractor and the details are given in the experiments. $\hat{\mu}$ and $\hat{\Sigma}$ could be estimated in the same way.
Our mapping $f$ is a pre-trained transformer encoder that uses a mask token prediction object. The loss is as follows:
(8) $\mathcal{L}_{\text{MTP}}=\frac{1}{k}\sum_{i}\frac{1}{m}\|s_{i}-e_{i}w\|^{2},$
where $k$ is the number of masked tokens, $s_{i}$ is the ground truth of $i$-th masked token, $e_{i}\!\in\!\mathbb{R}^{1\times d}$ denotes the features of time $i$ extracted by model, $d$ denotes the hidden dimension, and $w\in\mathbb{R}^{d\times m}$ is a linear projector.
3.4. ECG Model: Patch Step-by-Step Model
Overview. As shown in Figure 2, we introduce the Patch Step-by-Step Model (PSSM) for ECG, applying a hierarchical encoder-decoder architecture. The encoder stacks multiple patch operations and ConvBlock to progressively compress temporal resolution while doubling the hidden dimensions, effectively capturing multi-scale periodic patterns. The decoder reverses the patching process by progressively restoring the temporal resolution and halving the hidden dimensions by learnable unpatching operations.
Patching Encoder. For the encoder, the ECG signal $\bm{x}$ is first fed into the Embedding layer and then undergoes $l$ hierarchical patching operations, each followed by a ConvBlock. The computation formulations can be written as follows:
$\bm{e}^{1}&=\text{Embeding}(\bm{x}),\\ \bm{e}^{i+1}&=\text{ConvBlock}^{i}(\text{Patch}(\bm{e}^{i})),\\ \end{aligned}$
where the Embedding layer maps $\bm{x}\!\in\!\mathbb{R}^{t\times 1}$ to $\bm{e}^{1}\!\in\!\mathbb{R}^{t\times d}$, $d$ is the hidden dimension, $i=1,2,\dots,l$ denote the $i$-th layer of the encoder and $l$ denote the number of layers. The Patch layer splits $\bm{e}^{i}\in\mathbb{R}^{(t/2^{i-1})\times(2^{i-1}d)}$ into patches $\{(e^{i}_{2k-1}+e^{i}_{2k})/2\}$ with the length of $t/2^{i}$, and then fed to ConvBlocki to double the hidden dimension and derive $\bm{e}^{i+1}\in\mathbb{R}^{(t/2^{i})\times(2^{i}d)}$.
Unpatching Decoder. The encoder output is subsequently fed into the decoder, which performs $l$ unpatching operations before applying a final linear transformation to derive the ECG features. The formulations are:
$\bm{e}^{i+1}&=\text{ConvBlock}^{i}({\text{UnPatch}(\bm{e}^{i})}),\\ \bm{e}&=\bm{e}^{2l+1}w,\end{aligned}$
where $i\!=\!l\!+\!1,\dots,2l$ denotes $l$ decoder layers, $w$ is a linear layer and $\bm{e}$ denotes the final extracted features for ECG. The UnPatch layer restores $\bm{e}^{i+1}$ by weighted copying the features from the last layer and then fed into ConvBlock: $(e^{i+1}_{2k-1},e^{i+1}_{2k})$$=(\text{ConvBlock}^{i}(c_{1}e^{i}_{k}),$ $\text{ConvBlock}^{i}({c_{2}e^{i}_{k}}))$, where $c_{1}$ and $c_{2}$ are learnable parameters, and ConvBlock halves the hidden dimension.
Finally, after a linear projector $w$, the feature $\bm{e}$ of the ECG signal as the output of our PSSM will be further fed into downstream tasks described in Section 3.2.
Datasets Informer Medformer UniTS UniTS ECGPT Timer Timer PSSM
(Zhou et al., 2021) (Wang et al., 2024) (Gao et al., 2024) (Gao et al., 2024) (Davies et al., 2024) (Liu et al., 2024) (Liu et al., 2024) (Ours)
AF (Clifford et al., 2017) 0.648 0.970 0.653 0.944 0.952 0.996 0.999 0.999
NIFEADB (Behar et al., 2019) 0.753 0.876 0.720 0.756 0.750 0.887 0.936 0.937
CPSC2021 (Wang et al., 2021) 0.629 0.888 0.516 0.704 0.590 0.911 0.965 0.973
RFAA (Clifford et al., 2015) 0.535 0.659 0.414 0.579 0.528 0.875 0.943 0.937
SPB (Goldberger et al., 2000) 0.691 0.640 0.427 0.759 0.578 0.958 0.982 0.992
CPSC2018 (Liu et al., 2018) 0.858 0.563 0.277 0.463 0.383 0.630 0.768 0.843
Average 0.686 0.766 0.501 0.701 0.630 0.876 0.932 0.947
Table 1. Test Accuracy on ECG classification datasets. Bold numbers indicate the best performance, underlined numbers indicate the runner-up, and “*” denotes that the model is further pre-trained on the ECG datasets.
Dataset GQRS Informer Medformer UniTS UniTS ECGPT Timer Timer PSSM
(Xie et al., 2022) (Zhou et al., 2021) (Wang et al., 2024) (Gao et al., 2024) (Gao et al., 2024) (Davies et al., 2024) (Liu et al., 2024) (Liu et al., 2024) (Ours)
MITDB (Moody and Mark, 2001) 0.358 0.707 0.181 0.740 0.810 0.760 0.843 0.919 0.921
SVDB (Greenwald et al., 1990) 0.527 0.539 0.746 0.792 0.859 0.728 0.900 0.939 0.953
NFE (Silva et al., 2013a) 0.139 0.269 0.222 0.290 0.274 0.397 0.339 0.423 0.664
FEPL (Matonia et al., 2020) 0.246 0.848 0.577 0.730 0.701 0.672 0.658 0.803 0.978
CPSC2020 (Cai et al., 2020) 0.049 0.153 0.117 0.147 0.308 0.050 0.421 0.528 0.505
MITPDB (Maršánová et al., 2018) 0.090 0.119 0.115 0.346 0.619 0.364 0.667 0.791 0.901
Average 0.235 0.439 0.326 0.508 0.595 0.495 0.638 0.734 0.820
Table 2. Test F1 score on ECG detection datasets. Bold numbers indicate the best performance, underlined numbers indicate the runner-up, and “*” denotes that the model is further pre-trained on the ECG datasets. GQRS (Xie et al., 2022) is a traditional ECG method specifically designed for ECG detection, thus its results are reported only in this table.
4. Experiments
In this section, we first introduce the datasets for each task, then explain the implementation details, and finally analyze the proposed benchmark and method in detail.
4.1. Dataset
Building upon prior works (Davies et al., 2024; Sarkar and Etemad, 2021), we collected ECG evaluation datasets from: The China Physiological Signal Challenge 2018-2021 (CPSC2018, CPSC2019, CPSC2020, CPSC2021) (Liu et al., 2018; Gao et al., 2019; Cai et al., 2020; Wang et al., 2021), PhysioNet (Goldberger et al., 2000), as well as other ECG research datasets, with detailed descriptions following.
Classification. We collected 6 datasets for classification:
  • CPSC2018 contains 7 disease classes, including ST-segment depression, etc.
  • CPSC2021 is recorded from 12-lead Holter or 3-lead wearable ECG monitoring devices. Challenge focuses on atrial fibrillation diagnosis and has 3 classes in total.
  • AF Classification (AF) (Clifford et al., 2017) contains short-term ECG recordings with binary class for atrial fibrillation.
  • Non-Invasive Fetal ECG Arrhythmia (NIFEADB) (Behar et al., 2019) uses fetal ECGs for binary classification of arrhythmias.
  • Reducing False Arrhythmia Alarms (RFAA) (Clifford et al., 2015) designed to classify arrhythmia alarms in ICUs into 4 classes.
  • St. Petersburg Arrhythmia Database (SPB) (Goldberger et al., 2000) comprises 75 ECG recordings annotated with 6 diseases.
Detection. We evaluate 6 datasets for detection:
  • Fetal ECG with Pregnancy and Labor (FEPL) (Matonia et al., 2020) collects direct fetal ECGs and maternal abdominal ECGs annotated with fetal R-waves.
  • CPSC2020 was collected from arrhythmia patients with abnormal T-wave annotations.
  • MIT-BIH Arrhythmia Database (MITDB, MITPDB) (Moody and Mark, 2001; Maršánová et al., 2018), MITDB contains 48 ECG records and annotates the QRS complex, and MITPDB complements MITDB by adding P waves annotations.
  • Noninvasive Fetal ECG (NFE) (Silva et al., 2013a) collects maternal abdominal ECGs annotated with fetal R-waves.
  • Supraventricular Arrhythmia Database (SVDB) (Greenwald et al., 1990) contains 78 ECG records, annotated with QRS complexes.
Datasets Informer Medformer UniTS UniTS Timer Timer PSSM
(Zhou et al., 2021) (Wang et al., 2024) (Gao et al., 2024) (Gao et al., 2024) (Liu et al., 2024) (Liu et al., 2024) (Ours)
CPSC2019 (Gao et al., 2019) 2.241 1.866 4.172 1.975 1.988 0.839 0.480
DALIA (Reiss et al., 2019) 1.419 1.629 4.838 2.184 1.938 0.301 0.182
RDBH (Moody et al., 2014) 1.344 1.386 3.651 1.690 1.056 0.275 0.153
MIMICSub (Johnson et al., 2016) 0.538 0.384 4.306 2.093 1.091 0.100 0.117
NFE (Silva et al., 2013a) 1.400 1.064 2.196 1.089 1.289 0.353 0.311
PTB (Bousseljot et al., 1995) 0.618 0.301 1.491 0.667 0.480 0.065 0.025
Average 1.260 1.105 3.442 1.616 1.307 0.322 0.211
Table 3. Test FFD on ECG forecasting datasets. A smaller FFD is better. Bold numbers indicate the best performance, underlined numbers indicate the runner-up, and “*” indicates that the model is further pre-trained on the ECG datasets. ECGPT (Davies et al., 2024) is not applicable for this setting, and its performance is omitted.
Datasets Informer Medformer UniTS UniTS ECGPT Timer Timer PSSM
(Zhou et al., 2021) (Wang et al., 2024) (Gao et al., 2024) (Gao et al., 2024) (Davies et al., 2024) (Liu et al., 2024) (Liu et al., 2024) (Ours)
MITDB (Moody and Mark, 2001; Moody et al., 1984) 0.105 0.174 0.151 0.162 0.082 0.123 0.068 0.016
PTBXL (Wagner et al., 2020; Moody et al., 1984) 0.077 0.176 0.124 0.138 0.095 0.133 0.054 0.019
ADFECGDB (Jezewski et al., 2012) 0.495 0.347 0.430 0.417 0.457 0.323 0.321 0.249
FEPL (Matonia et al., 2020) 0.982 1.071 0.911 1.304 1.219 1.344 0.511 0.081
BIDMC (Pimentel et al., 2016) 1.562 1.049 1.033 1.275 1.506 1.026 0.959 0.218
SST (Lazović et al., 2024) 0.772 0.793 0.747 0.731 0.728 0.699 0.526 0.216
Average 0.665 0.601 0.566 0.671 0.681 0.608 0.407 0.133
Table 4. Test FFD on ECG generation datasets. A smaller FFD is better. Bold numbers indicate the best performance, underlined numbers indicate the runner-up, and “*” indicates that the model is further pre-trained on the ECG datasets.
Forecasting. While any ECG datasets can be used for forecasting, for broader validation, we selected 6 datasets designed for different research objectives:
  • CPSC2019 (Gao et al., 2019) includes 2,000 single-lead ECG recordings collected from patients with cardiovascular disease. It is often used in ECG detection tasks.
  • PPG-DaLiA (DALIA) (Reiss et al., 2019) contains Contains ECG, photoplethysmography, and 3D accelerometer data. It is often used in ECG generation tasks.
  • PTB Diagnostic ECG Database (PTB, PTBXL) (Bousseljot et al., 1995; Wagner et al., 2020), PTB contains 549 records from 290 subjects. PTBXL ECG contains 21,799 clinical 12-lead ECGs from 18,869 patients, with each ECG lasting 10 seconds. They are often used in ECG classification tasks.
  • A subset of MIMIC-III Waveform Database Matched Subset (MIMICSub) (Johnson et al., 2016), the original dataset contains 22,247 numerical records of 10,282 different ICU patients. We selected a portion of this dataset.
  • Detection of Heart Beats Dataset (RDBH) (Moody et al., 2014) contains 10-minute recordings. Each recording contains four to eight signals. It is often used in ECG detection tasks.
  • European ST-T Database (EBD) (Taddei et al., 1992) contains 90 annotated Holter recordings from 79 subjects. It is often used in ECG detection tasks.
Generation. Generation evaluation contains 6 datasets, including FEPL. Like previous work(Chatterjee et al., 2020), we augment the MITDB and PTBXL(Wagner et al., 2020) by injecting noise from the MIT-BIH Noise Stress Test Database (Moody et al., 1984) to build denoising datasets. The remaining datasets are described below:
  • Abdominal and Direct Fetal ECG (ADFECGDB) (Jezewski et al., 2012) contains maternal abdominal ECG and direct fetal ECG obtained from 5 pregnant subjects.
  • BIDMC PPG and Respiration Dataset (BIDMC) (Pimentel et al., 2016) extracts ECG from the MIMIC II Waveform Database (Saeed et al., 2011), with annotated photoplethysmogram signals.
  • SensSmartTech Database (SST) (Lazović et al., 2024) includes synchronized cardiovascular signals from ECG, phonocardiograms, and photoplethysmography.
We further organized an extension ECG Dataset for LTMs pre-training, comprising AF, CPSC2018, SPB, PTB, PTBXL, NFE, MIMICSub, The Georgia 12-lead ECG Challenge Database (Goldberger et al., 2000), and the Shaoxing and Ningbo Hospital ECG Database (Zheng et al., 2020b, a). This dataset contains 98,513 subjects processed into 8,414,834 training samples. Although the extension ECG Dataset overlaps with other datasets, the training objectives differ (e.g., AF and CPSC2018), or the overlapping samples are removed (e.g., MIMICSub). We describe more details about ECG datasets in the supplementary material.
Figure 3. Qualitative results of different models on the ECG generation task. The three rows of samples were selected from the MITDB (Moody and Mark, 2001), FEPL (Matonia et al., 2020), and SST (Lazović et al., 2024) datasets, respectively. Red curves are the ground truth ECG, and blue curves are the generated ECG.
[…]
4.2. Implementation Details
Datasets. All ECG recordings were resampled to 100Hz with 500 points fixed length. We discarded recordings with $>25\%$ missing data or abnormal morphologies, imputing missing segments via linear interpolation. Multi-channel ECG recordings were split into single-channel recordings to standardize input dimensions. The datasets were partitioned into 50% for training/fine-tuning and 50% for testing. All datasets have been stripped of sensitive information.
We addressed class imbalance by removing underrepresented classes and downsampling overrepresented ones for classification. In detection tasks with severe class imbalance (waveform annotation positions <5%), we adopted the F1-score. Non-maximum suppression was applied to eliminate redundant waveform probabilities. The predicted position must be within a $\pm$70ms window centred on the annotation position(Cai et al., 2020; Association for the Advancement of Medical Instrumentation et al., 1998). For forecasting, the predicting length is 100.
Models. We evaluated advanced LTMs introduced in Section 2: Timer (Liu et al., 2024), UniTS (Gao et al., 2024), and ECGPT (Davies et al., 2024). Both raw versions and variants are pre-trained on the extension ECG Dataset, denoted by “*”. The mapping $f$ of FFD is the LTM constructed by the transformer encoder. Additionally, we tested state-of-the-art transformer-based end-to-end models: Informer (Zhou et al., 2021) and Medformer (Wang et al., 2024) alongside our proposed PSSM.
All pre-trainings use the extension ECG Dataset and were conducted on NVIDIA RTX 4090 GPUs with a batch size of 8192 over 300 epochs. We use the Adam optimizer (Kingma and Ba, 2014) for model optimization with a learning rate of $1\times 10^{-6}$. All test experiments were conducted on NVIDIA A800 GPUs with a batch size of 1024 over 100 epochs, using the Adam optimizer with learning rates empirically selected from $\{1\!\times\!10^{-4},5\!\times\!10^{-5},1\!\times\!10^{-5}\}$. For LTMs, we freeze the backbone and only fine-tune a fully connected layer and task heads, yet remain unmodified in forecasting. Other models undergo full training. More details are provided in the supplementary material.
4.3. Analysis
Through the following Q&A, we provide an in-depth analysis of the experimental results, focusing on the rationale of our ECG benchmark, the robustness of our proposed FFD, and the effectiveness of our PSSM.
Q1. Is the proposed ECG benchmark reasonable and is the PSSM effective?
A1.
The main results for our four evaluation tasks, i.e., classification, detection, forecasting, and generation, are presented in Table 1 to Table 4, respectively. Our proposed method, PSSM, achieves state-of-the-art performance across all tasks, with an average performance of 0.947 in classification accuracy, 0.820 in detection F1 score, 0.211 in forecasting FFD, and 0.133 in generation FFD.
Compared to traditional methods such as Medformer, PSSM achieves an averaged relative improvement of 83.4%, specifically 23.6% in classification accuracy, 151.2% in detection F1 score, 80.9% in forecasting FFD, and 80.9% in generation FFD. Compared to large time-series methods such as Timer, PSSM achieves an averaged relative improvement of 49.6%, specifically 8.1% in classification accuracy, 28.6% in detection F1 score, 83.8% in forecasting FFD, and 78.1% in generation FFD. Compared to rule-based GQRS in detection, PSSM improves significantly by 151.2%. These results demonstrate the effectiveness of our method and highlight the need to develop specialized models for ECG analysis rather than relying solely on large time-series models.
These experimental results further validate the robustness and consistency of our proposed ECG benchmark. For example, our method PSSM consistently achieves the highest average performance, while Timer consistently secures the runner-up position, demonstrating the consistency of four evaluations. Even when some models perform well on specific datasets (e.g., Informer on CPSC2018 classification), our comprehensive benchmark effectively mitigates these outlier cases, demonstrating robustness.
Figure 4. Comparisons for MSE and FFD under different temporal shifts. The top row displays ECG corresponding to various temporal shifts, where red curves are the ground truth and blue curves are the ECG with time shift. The bottom row illustrates how MSE and FFD vary across different temporal shifts. As temporal shift increases, although the semantic meaning of ECG remains consistent, the MSE increases while the FFD remains stable.
[…]
Q2. Can FFD effectively assess ECG quality?
A2.
Our proposed FFD is a reliable metric for assessing the quality of generated ECGs, with lower FFD values indicating better ECG fidelity. For example, as illustrated in the third row of Figure 3, PSSM accurately reconstructs the R-wave locations and amplitudes from phonocardiograms on the SST dataset, whereas other models yield clinically meaningless flat lines. This finding aligns with the quantitative results in Table 4, where PSSM achieves the best FFD as 0.216, significantly outperforming other methods, such as Timer* by 0.31 (59%). This consistency between qualitative and quantitative results is also validated in the first-row MITDB and second-row FEPL datasets.
Additionally, FFD shows robustness in evaluating the semantics of generated ECGs. As illustrated in Figure 4, ECGs preserve diagnostic semantics under temporal shifts in the top row, with FFD remaining stable at 0.022 despite perturbations, as shown in the bottom. In contrast, the MSE sharply increases from 0.281 to 1.579 under the same shifts, failing to capture the unchanged semantics of ECG signals. This demonstrates that FFD is robust to temporal fluctuations, only related to ECG semantics.
In conclusion, FFD is an effective and robust metric for assessing the quality of ECGs. For completeness, in the supplementary materials, we supplemented the MSE performance in the forecasting and generation tasks, where our PSSM is still the best model, and we also provided more detailed discussions on MSE and FFD.
Q.3 How does hierarchical patching improve PSSM?
A.3
To validate the hierarchical patching strategy, as shown in Table 5, we conducted two ablation studies: (1) replacing ConvBlocks with a transformer encoder (“rp trans”) and (2) applying patching only at the input stage, similar to LTMs, instead of using hierarchical patching (“w/o ssp”). The results indicate that PSSM with ConvBlocks and hierarchical patching achieves the best performance, outperforming the “rp trans” variant by 63.2%, 159.8%, 85.6%, and 87.9% across the four evaluation tasks, respectively, and outperforming the “w/o ssp” variant by 9.6%, 70.06%, 78.3%, and 62.4% across the four evaluation tasks. These results demonstrate that the hierarchical patching strategy is both necessary and effective. The possible reason for the failure of the transformer encoder is that hierarchical patching could capture the quasi-periodic temporal patterns for accurate feature extraction, while “rp trans” variants create complex feature interactions by self-attention along the time dimension that are redundant for periodic signals.
Model CLS DET FCAST GEN
(Accuracy) (F1) (FFD) (FFD)
   rp trans 0.582 0.316 1.198 1.101
   w/o ssp 0.866 0.483 0.791 0.353
PSSM 0.949 0.820 0.172 0.133
Table 5. Ablation studies of our PSSM on four ECG evaluations, where CLS, DET, FCAST, and GEN denote classification, detection, forecasting, and generation, respectively. “rp trans” substitutes ConvBlocks with Transformer encoders, and “w/o ssp” has no hierarchical patch.
Figure 5. Illustrations of the attention maps for a pre-trained time-series model with and without additional ECG pre-training. Timer, which is further pre-trained on ECG datasets, exhibits an attention map with periodic patterns not observed in the raw Timer model. The samples are selected from the PTB dataset (Bousseljot et al., 1995).
[…]
Q.4 Are LTMs suitable for ECG tasks?
A.4
Based on our experiments, the performance of LTMs is less than ideal. As shown in Table 1 to Table 4 and Figure 3, LTMs, including Timer, Timer, UniTS, UniTS, and ECGPT, exhibit suboptimal performance in our ECG benchmark. Specifically, (a) the raw LTMs (i.e., Timer and UniTS) significantly underperform (e.g., UniTS and Timer show accuracy drops of 0.446 and 0.071 compared to PSSM in classification). (b) ECGPT, an LTM specifically pre-trained on ECG datasets, also struggles (e.g., with a 39.6% lower F1-score than PSSM in detection). (c) The LTMs further pre-trained on ECG datasets, e.g., Timer outperform Timer by 75.3% in forecasting FFD yield 34.5% lower than PSSM. Demonstrating some improvement compared to their raw versions, but still do not match the performance of our proposed model.
On the other hand, the qualitative results show that raw LTMs fail to capture the periodic nature of ECG signals. Figure 5 reveals that the raw Timer model attends only to nearby tokens, whereas Timer (further pre-trained on ECG data) adjusts its attention to capture periodic patterns. For example, Timer focuses on even tokens in Example A and alternates between even and odd tokens in Example B. This phenomenon indicates that the underlying attention mechanism of transformers in current LTMs is unsuitable for ECG data. Even when further pre-trained on ECG datasets to learn more ECG patterns like Timer, transformers still fail to fully leverage their potential (i.e., half of the attention weights are wasted), resulting in suboptimal performance. Therefore, an improved model architecture is necessary to achieve better performance for ECG analysis. In the Supplementary Material, we also provide further evidence and discussion of the unsuitability of LTMs for ECG tasks.
5. Conclusion
This paper introduces a comprehensive benchmark for ECG analysis, including classification, detection, forecasting, and generation, designed to address real-world clinical needs for ECG, such as disease diagnosis and early risk alerts. Second, we propose a novel metric for ECG, Feature-based Fréchet Distance (FFD), which addresses the limitations of Mean Squared Error (MSE) in ECG assessment, particularly in capturing the semantic fidelity of ECG. Additionally, we introduce the Patch Step-by-Step Model (PSSM), an ECG-specialized architecture that employs a hierarchical patching strategy to effectively capture the nature of ECG signals. Extensive experiments validate the rationality of our ECG benchmark, the robustness of our proposed FFD, and the superiority of our PSSM compared to other state-of-the-art time-series methods across all evaluation tasks. Our future work focuses on developing ECG-optimized time-series model pre-training to bridge the gap between general LTM pretraining and ECG requirements.